If it requires 75.0 mL of 0.500M NaOH to neutralize 165.0 mL of an HCI solution, what is the concentration of the HCI solution?


1
Expert's answer
2021-12-02T01:03:16-0500

Solution:

Balanced chemical equation:

NaOH + HCl → NaCl + H2O

According to stoichiometry:

n(NaOH) = n(HCl)

CM(NaOH) × V(NaOH) = CM(HCl) × V(HCl)

Therefore,

CM(HCl) = CM(NaOH) × V(NaOH) / V(HCl)

CM(HCl) = (0.500 M × 75.0 mL) / (165.0 mL) = 0.227 M

CM(HCl) = 0.227 M


Answer: The concentration of the HCI solution is 0.227 M